OpenStudy (anonymous):

1. . Evaluate: -6 - (-1) -7 5 -5 6 2. Evaluate: 32 + (6 - 2) • 4 - 6/3 20 23 50 26 3. Evaluate: [(40 + 5) - 32] ÷ 9 • 2 8.6 43 8 2

OpenStudy (dobby1):

the first one is -5

OpenStudy (anonymous):

@Dobby1 Thanks!

OpenStudy (anonymous):

do you happen to know # 2 and #3?

OpenStudy (anonymous):

1)we have : -6+1 = -5 2)we should do : (6 - 2) first : 4 then 4*4=16 then have : 32+8-6/3=40-6/3 you do next ! 3)45-32=13 then : 9^2=81 Then : 13/81

OpenStudy (dobby1):

for number 2 this is what your question looks like after you simplify the fraction and parenthases 32 + 4 * 4 - 2

OpenStudy (dobby1):

so after you multiply you get 32+16-2

OpenStudy (dobby1):

after you add you get 48-2

OpenStudy (dobby1):

when you subtract you get 46

OpenStudy (dobby1):

but that is not an answer so i would say 50 i gues

OpenStudy (anonymous):

@Dobby1 Thanks!

OpenStudy (dobby1):

what is the 2 in the third one?

OpenStudy (anonymous):

Idk thats alll that the question said

OpenStudy (anonymous):

i keep getting 46 for #2 and #3 is 2 i guess